40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7100     Entry time: Tue Aug 7 03:20:56 2012
Author: Jenne 
Type: Update 
Category: ASS 
Subject: ASS setup, on, off scripts written 

I wrote new setup, on and off scripts for the arm ass.  They take the arm as an argument, so it's the same script for both arms.  Scripts are in ...../scripts/ASS/ , and have been checked in to the 40m svn.

So far the on script doesn't really do anything, since I haven't chosen values for the CLKGAINs of the lockins.  The old values were 30 for lockins 12, 14, 27, 29 and 250 for lockins 7, 9, 22, 24.  Unfortunately, I have no memory of which lockin means what in the old numbered system.  I'll have to look that up somehow.  Or, just dither the optics using some value and look at the spectrum to see the resulting SNR and just pick something that gives me reasonable SNR.

I modified the ASS model slightly:

* Added an overall gain to the ASS_DOF2 library part, between the matrix and the servo inputs so we can do soft startups.  Self - remember that the main ASS screen needs to be modified to reflect this!

* Rearranged the order that the demodulated signals go into the matrix.  I hadn't paid attention, and the old ordering had the transmission (TRX/TRY) demod signals interleaved with the LSC demod signals.  I've changed it to be all the TR signals, then all the LSC signals.  I think this makes more sense, since we will use these inputs separately, so now they're on different halves of the matrix.

ELOG V3.1.3-